# Hybrid optimal control with mixed-integer Lagrangian methods

### v1

@@ 5 Conclusions
This work presents an efficient method to address hybrid optimal control problems without any relaxations. The proposed approach is based on the augmented Lagrangian framework and a routine for nonlinear optimization with mixed-integer linear constraints. The robustness and performances of the algorithm were assessed on three problem examples from different fields of application. Promising preliminary results validate the strategy here developed.
Future research should consider the convergence analysis of Algorithm 3.1 and the possibility to integrate other techniques for handling nonlinear constraints.

### v2

@@ 5 Conclusions
This work presents an efficient method to address hybrid optimal control problems without any integrality relaxations. The proposed approach is based on the augmented Lagrangian framework combined with a routine for nonlinear optimization with mixed-integer linear constraints. The robustness and performance of the algorithm were assessed on two problem examples from different fields of application, validating the proposed strategy. Future research should consider the integration of other techniques for handling nonlinear constraints.
Acknowledgements
We thank Rebecca Richter for providing the dynamic programming model and implementation.